# Slice-100K: A Multimodal Dataset for Extrusion-based 3D Printing

### v2

@@ 3.2   G-code Translation
G-code translation involves translating a G-code from one flavor to another while preserving the necessary context associated with each flavor and finding a correspondence between any two given flavors of G-code. We begin with G-code data in two different flavors, Sailfish and Marlin. Sailfish is a legacy G-code format that is not currently used by the 3D printing community. Marlin is a modern G-code format that has been heavily adopted, and in some cases, other G-code flavors are built on top of Marlin. Given this, we leverage our proposed dataset to finetune GPT-2 for the task of G-code translation. G-code is inherently a low-level language, and for a task like translation, the quality of data being fed into an LLM has a significant impact on its performance. Keeping this in mind, we perform some data pre-processing to effectively maintain the context across lines of G-code.
@@ 3.2.1   Data Pre-processing Methods
A major challenge in applying language-modelling based techniques to G-code is the length of G-code files. While a shape’s G-code representation can be separated into layers (which do not share information and can therefore be handled independently) this is still not sufficient as a single layer from our dataset can be over the token limit. This motivates methods for further splitting of G-code layers, allowing us to decompose mappings between G-code files into a series of mappings between smaller G-code portions.
Crucially, these methods can be applied to different G-codes regardless of the variants they are written in while ensuring that the resulting pairs of G-code segments represent the same spatial semantics. To accomplish this, we first permute the contours in each G-code layer so that they have the same ordering. Following that, we adaptively select portions to create matching pairs.
@@
Contour Flipping: Let $L_{A}$ and $L_{B}$ be two layers G-code layers which use different flavors to represent the same 3D information. We can decompose each of these layers into a series of $N$ contours $c^{(A)}_{1},..,c^{(A)}_{N}$ and $c^{(B)}_{1},..,c^{(B)}_{N}$, each represented using their respective flavor. Disregarding the difference in flavors, both sequences contain the same set of unique of contours. Because of this, we can define a bijective mapping $M:[N]\rightarrow[N]$ such that the contour $c^{(B)}_{M(i)}$ is equivalent to $c^{(A)}_{i}$.
@@
The primary challenge in our preprocessing is to find this bijection, which once found allows us to re-order the contours of $L_{B}$ so that $\forall i\in[N]$ we have that $c^{(B)}_{i}$ is equivalent to $c^{(B)}_{i}$. To determine $M$, we iterate over each contour $c^{(A)}_{i}$ in $L_{A}$ and find its corresponding contour $c^{(B)}_{M(i)}$ in $L_{B}$. We consider two contours to be matching if there are specific commands which are included in both. More specifically, we define a method for representing a single line of G-code so that identical representations will indicate matching contours.
@@
To minimize the possibility of a duplicate representation (that could lead to a false match), we base this criteria on G-code lines which contain commands to extrude at specified coordinates. Other commands are disregarded as they are likely to be repeated throughout a file or contain syntax which differs across flavors. In contrast, extrusion locations are specified using floating point coordinates with several digits of precision making it rare for the same point to appear in different contours. We further account for the possibility of duplicate locations by concatenating the line’s coordinates with the next two lines in the contour where possible. If these following lines do not contain a location-specific extrusion command, we simply include in their place a token denoting an empty line. Together, this creates a string representation of each line that strips away flavor-specific syntax while including enough contextual information to prevent unwanted duplicates.
@@
Using this consistent characterization of G-code lines allows us to match contours by simply finding a single pair of lines with the same representation. However, due to the length of G-code layers, it is highly inefficient to consider all possible pairs of lines when looking to match contours. To alleviate this, we pre-compute a lookup table for $L_{B}$. For each line of a contour $c^{(i)}_{B}$ the lookup table maps from the line representation to the index $i$. Then, when iterating over the contours of $L_{A}$ we simply compute the representation for each line and search the lookup table. If there is a match, then we add these indices to our bijection $M$. While this contour flipping method cannot be guaranteed to always find the correct bijection $M$ due to variations amongst some contours, we find that it is highly reliable, producing aligned G-code for over $99.9\%$ of the G-code layers in our dataset. We include pseudocode for our method in the Appendix (Algorithm 1).
@@
Pair Creation: Given two G-code layers which have undergone contour flipping so that they have the same high-level semantic ordering, we can reasonably expect to divide them each into pairs of contiguous sections sharing the same 3D information. Because there are often commands included in one flavor but not the other, we cannot simply select portions of equal length and expect them to be translatable. Instead, we have to adaptively determine the cutoff points for each section.
@@
Our pair creation approach finds these matching line indices while respecting a maximum length parameter (see Algorithm 2 in Appendix). In short, we iteratively find index $k^{A}_{i+1}$ by starting with a candidate value which is $k^{A}_{i+1}$ plus the maximum length. We then try to find a matching line in $L_{B}$ and, if successful, consider this a pair. If we cannot find a matching line for our candidate, we decrement the candidate line index by one and continue trying. We use a line representation similar to the one used for contour flipping to determine whether a pair of lines are matching.
Handling Extrusion Values: Through the previously described preprocessing methods we have been able to create pairs of G-Code chunks which represent the same local information and can therefore be translated between. However, there is an additional non-local dependence which must be accounted for in the G-Code extrusion values. In addition to telling the 3D printer where to move, a line of G-Code also tells it how much material to extrude during this movement.
This is specified through an "E" command which states how much total material will have been extruded once that point is reached. For instance, if one line contains an E value of $3$ and the next line has an E value of $3.1$, then .1.1.1.1 units of material should be extruded during this movement. There are also specialized language-specific commands throughout a shape’s G-Code which reset the total extrusion values to some smaller constant.
Because these values represent a cumulative sum of all material extruded up to that point starting from the most recent reset value, there is a non-locality element that must be addressed. During preprocessing, we amend each extrusion value by subtracting the previous line’s extrusion value. We call this new value the relative extrusion. This represents only the amount of material that is to be extruded during this movement and allows for any translation model to learn a simple local mapping that is not dependent on other chunks. Finally, after generating G-Code in this relative form, we convert it back to its original format by computing its cumulative sum.
@@ 3.3   Geometric Transformation - Scaling
At the time of our evaluation, we empirically arrived at the maximum number of lines of G-code an LLM in our test suite can accept before crossing their respective token limits. We leverage this fact to chunk the G-code.
@@ 3.4   Evaluation Metrics
To measure the quality of G-code generation models, we introduce an image-space IoU metric for G-code fidelity in comparison to a ground truth. Because small errors in the produced G-code can lead to significant divergence in the produced shape, we find it insufficient to use a language-based metric for evaluation. Instead, we use an image-based measure of fidelity by rendering top-down images of each layer.
G-code Renderer: To the best of our knowledge, there does not currently exist an open-source software which can programatically generate renderings of G-code objects. To remedy this, we introduce and make public our Python-based tool for this purpose. The renderer generates a top-down rendering of an individual layer. This layer-wise approach is sensible for examining the 3D structure as it avoids occlusions and captures all relevant parts of the shape, even the infill that provides internal structural support for the part and may not be visible from the outside.
@@ 4   Experiments
We use Slice-100K for two tasks: evaluating current LLMs for G-code geometric transformation (scaling) and G-code flavor translation by finetuning GPT2,
@@ 4.1   Evaluating Existing LLMs for G-Code Geometric Transformations
We evaluate some of the existing chat-based LLMs (GPT series [58, 59], Claude [60], Llama-2 [42], and Starcoder [61]) for performing geometric transformations, specifically scaling a layer of the model. We find GPT-3.5 and GPT-4 struggle with the S-shape. Claude-2 is able to generate the outer contour of the cylinder and cube but struggles with infill generation for the cylinder and the S-shape. Furthermore, we see that the open-source models—Llama-2-70b and Starcoder—do not perform well. We visualize the G-code outputs from the various LLMs in our test suite and render them using Ultimaker’s Cura [62] slicing software. Our results are shown in Figure 4.
{tblr}
colspec = X[0.18,l,t]X[0.18,c,t]X[0.18,c,t]X[0.18,c,t], stretch = 0, rowsep = 4pt, hlines = black, 1pt, vlines = black, 1pt, Expected & Refer to caption Refer to caption Refer to caption
@@ 4.2   G-code Flavor Translation
For finetuning, we create a paired dataset of G-Code chunks in each flavor using the preprocessing methods outlined in Section 3.2.1 using a maximum chunk size of $20$ lines. We then finetune a lightweight GPT-2 model for translation using a next-token prediction loss. During inference, we do not have access to the ground-truth Marlin G-code which would be needed to determine the cutoff lines for pair creation, so we instead split our Sailfish input into smaller chunks of fixed-length.
@@
Table 3 shows the results of performing finetuning on differently sized subsets of Slice-100K. We denote a GPT-2 model finetuned on $n$ shapes from our dataset as GPT-$2^{(n)}$. Figure 5 shows example renderings of shapes that have undergone flavor translation by our GPT-25 model. We also include an example G-Code that has been translated in the Appendix (Figure 9). We find that even finetuning on very small subsets of our dataset leads to greatly enhanced G-Code translation abilities. Increasing the amount of training data beyond just five G-Code shapes finetuning ceases to yield improvements. We attribute this to our preprocessing methods which reduce the complex translation task to a simple local mapping, thereby reducing the amount of data needed for learning.
@@ 5   Conclusion
In this paper, we presented Slice-100K, the first large-scale, curated dataset of over 100,000 G-code files along with their corresponding STL CAD files, renderings, and metadata. This dataset addresses a significant gap in the availability of comprehensive resources for additive manufacturing. Our evaluation demonstrated the utility of Slice-100K in utilizing existing language models for tasks such as G-code debugging, geometric transformations, and comprehension. Additionally, we introduced a novel application of using Slice-100K for LLM-based G-code flavor translation, showcasing the potential of our dataset in advancing the field. We believe that Slice-100K will serve as a foundational resource for future innovations in manufacturing, paving the way for the development of domain-specific foundation models.
Limitations: Despite these advancements, Slice-100K has certain limitations. One major challenge is the difficulty in verifying the LVIS categories for the models. Additionally, all models in Slice-100K were sliced only along the default Z-direction. This uniform slicing approach may limit the dataset’s applicability for research into multi-directional slicing techniques and their impact on manufacturing outcomes. Addressing these limitations in future versions of the dataset will be crucial for further enhancing its utility and broadening its application scope.
@@ Acknowledgements
This work was supported by the National Science Foundation under grant CMMI-AM-2347623/2347624. We would like to thank the NVIDIA Corporation for providing GPUs for academic research.
@@ Appendix A Appendix
4:     $Lookup\leftarrow HashMap()$triangleright Create hash index of contours in Layer B
5:     for $c\leftarrow 1$ to length($ContoursB$do
6:         for $i\leftarrow 1$ to length($ContoursB[c]$do
@@
15:     $Mapping\leftarrow HashMap()$triangleright Find bijection between layers
16:     for $c_{A}\leftarrow 1$ to length($ContoursA$do
@@
26:     for $i\leftarrow 1$ to length($ContoursB$do
@@
14:         if $found$ then $\triangleright$ Add matching pair of chunks to dataset
@@
19:              $end_{i}=end_{i}-1$triangleright Could not find a line matching line $end_{i}$, try a smaller chunk

### v3

@@ 3.2 G-code Translation
G-code translation involves converting a G-code from one flavor to another while preserving the necessary context associated with each flavor and finding a correspondence between any two given flavors of G-code. We begin with G-code data in two different flavors, Sailfish and Marlin. Sailfish is a legacy G-code format that is not currently used by the 3D printing community. Marlin is a modern G-code format that has been heavily adopted, and in some cases, other G-code flavors are built on top of Marlin. Given this, we leverage our proposed dataset to finetune GPT-2 for the task of G-code translation from Sailfish to Marlin. G-code is inherently a low-level language, and for a task like translation, the quality of data being fed into an LLM has a significant impact on its performance. Keeping this in mind, we perform some data pre-processing to effectively maintain the context across lines of G-code.
@@ 3.2.1 Data Pre-Processing Methods
A major challenge in applying language-modeling-based techniques to G-code is the length of G-code files. While the G-code representation of a 3D shape can be separated into layers (which do not share information and hence be handled independently), this is still not sufficient, as a single layer can be over the token limit. This motivates methods for further splitting of G-code layers, allowing us to decompose mappings between G-code files into a series of mappings between smaller G-code portions. Crucially, these methods can be applied to different G-codes regardless of the variants they are written in while ensuring that the resulting pairs of G-code segments represent the same spatial semantics. To accomplish this, we first permute the contours in each G-code layer so that they have the same ordering. We then adaptively select portions to create matching pairs.
@@
Contour Flipping: Let $L_{A}$ and $L_{B}$ be two G-code layers that use different flavors to represent the same 3D information. We can decompose each of these layers into a series of $N$ contours $c^{(A)}_{1},..,c^{(A)}_{N}$ and $c^{(B)}_{1},..,c^{(B)}_{N}$, each represented using their respective flavor. Both sequences contain the same set of unique contours irrespective of the flavors. Because of this, we can define a bijective mapping $M:[N]\rightarrow[N]$ such that the contour $c^{(B)}_{M(i)}$ is equivalent to $c^{(A)}_{i}$.
@@
The primary challenge in our preprocessing is to find this bijection, which, once found, allows us to re-order the contours of $L_{B}$ so that $\forall i\in[N]$ we have that $c^{(B)}_{i}$ is equivalent to $c^{(B)}_{i}$. To determine $M$, we iterate over each contour $c^{(A)}_{i}$ in $L_{A}$ and find its corresponding contour $c^{(B)}_{M(i)}$ in $L_{B}$. We consider two contours to be matching if there are specific commands which are included in both. More specifically, we define a method for representing a single line of G-code so that identical representations will indicate matching contours.
@@
To minimize the possibility of a duplicate representation (that could lead to a false match), we base this criterion on G-code lines, which contain commands to extrude at specified coordinates. Other commands are disregarded as they are likely to be repeated throughout a file or contain syntax that differs across flavors. In contrast, extrusion locations are specified using floating point coordinates with several digits of precision, making it rare for the same point to appear in different contours. We further account for the possibility of duplicate locations by concatenating the line’s coordinates with the next two lines in the contour where possible. If these following lines do not contain a location-specific extrusion command, we simply include in their place a token denoting an empty line. Together, this creates a string representation of each line that strips away flavor-specific syntax while including enough contextual information to prevent unwanted duplicates.
@@
Using this consistent characterization of G-code lines allows us to match contours by simply finding a single pair of lines with the same representation. However, due to the length of G-code layers, it is highly inefficient to consider all possible pairs of lines when looking to match contours. To alleviate this, we pre-compute a lookup table for $L_{B}$. For each line of a contour $c^{(i)}_{B}$, the lookup table maps from the line representation to the index $i$. Then, when iterating over the contours of $L_{A}$, we compute the representation for each line and search the lookup table. If there is a match, then we add these indices to our bijection $M$. While this contour flipping method cannot be guaranteed to always find the correct bijection $M$ due to variations amongst some contours, we find that it is highly reliable, producing aligned G-code for over $99.9\%$ of the G-code layers in our dataset. We include pseudocode for our method in the Appendix (Algorithm 1).
@@
Pair Creation: Given two G-code layers that have undergone contour flipping so that they have the same high-level semantic ordering, we can reasonably expect to divide them each into pairs of contiguous sections sharing the same 3D information. Because there are often commands included in one flavor but not the other, we cannot simply select portions of equal length and expect them to be translatable. Instead, we have to adaptively determine the cutoff points for each section.
@@
Our pair creation approach finds these matching line indices while respecting a maximum length parameter (see Algorithm 2 in Appendix). In short, we iteratively find index $k^{A}_{i+1}$ by starting with a candidate value which is $k^{A}_{i+1}$ plus the maximum length. We then try to find a matching line in $L_{B}$ and, if successful, consider this a pair. If we cannot find a matching line for our candidate, we decrease the candidate line index by one and continue trying. We use a line representation similar to the one used for contour flipping to determine whether a pair of lines is matching.
Handling Extrusion Values: Through the previously described preprocessing methods we have been able to create pairs of G-code chunks which represent the same local information and can therefore be translated between. However, there is an additional non-local dependence that must be accounted for in the G-code extrusion values. In addition to telling the 3D printer where to move, a line of G-code also tells it how much material to extrude during this movement.
This is specified through an "E" command which states how much total material will have been extruded once that point is reached. For instance, if one line contains an E value of $3$ and the next line has an E value of $3.1$, then .1.1 units of material should be extruded during this movement. There are also specialized language-specific commands throughout a shape’s G-code, which reset the total extrusion values to some smaller constant.
Because these values represent a cumulative sum of all material extruded up to that point starting from the most recent reset value, there is a non-locality element that must be addressed. During preprocessing, we amend each extrusion value by subtracting the previous line’s extrusion value. We call this new value the relative extrusion. This represents only the amount of material that is to be extruded during this movement and allows for any translation model to learn a simple local mapping that is not dependent on other chunks. Finally, after generating the G-code in this relative form, we convert it back to its original format by computing its cumulative sum.
@@ 3.3 Geometric Transformation - Scaling
At the time of our evaluation, we empirically arrived at the maximum number of lines of G-code an LLM in our test suite can accept before crossing their respective token limits. We leverage this fact to chunk the G-code before feeding it to an LLM.
@@ 3.4 Evaluation Metrics
To measure the quality of G-code generation models, we introduce an image-space IoU metric for G-code fidelity in comparison to ground truth. Because small errors in the produced G-code can lead to significant divergence in the produced shape, we find it insufficient to use a language-based metric for evaluation. Instead, we use an image-based measure of fidelity by rendering top-down images of each layer.
G-code Renderer: To the best of our knowledge, there does not currently exist open-source software that can programmatically generate renderings of G-code objects. To remedy this, we introduce and make public our Python-based tool for this purpose. The renderer generates a top-down rendering of an individual layer. This layer-wise approach is sensible for examining the 3D structure as it avoids occlusions and captures all relevant parts of the shape, even the infill that provides internal structural support for the part and may not be visible from the outside.
@@ 4 Experiments
We use Slice-100K for two tasks: evaluating current LLMs for G-code geometric transformation (scaling) and G-code flavor translation by finetuning GPT-2,
@@ 4.1 Evaluating Existing LLMs for G-code Geometric Transformations
We evaluate some of the existing chat-based LLMs (GPT series (Brown et al., 2020; OpenAI, 2023), Claude (Anthropic, 2023), Llama-2 (Touvron et al., 2023), and Starcoder (Li et al., 2023b)) for performing geometric transformations, specifically scaling a layer of the model. We find GPT-3.5 and GPT-4 struggle with the S-shape. Claude-2 is able to generate the outer contour of the cylinder and cube but struggles with infill generation for the cylinder and the S-shape. Furthermore, we see that the open-source models—Llama-2-70b and Starcoder—do not perform well. We visualize the G-code outputs from the various LLMs in our test suite and render them using Ultimaker’s Cura (Ultimaker, 2023) slicing software. Our results are shown in Figure 4.
Expected Refer to caption Refer to caption Refer to caption
@@ 4.2 G-code Flavor Translation
For finetuning, we create a paired dataset of G-code chunks in each flavor using the preprocessing methods outlined in Section 3.2.1 using a maximum chunk size of $20$ lines. We then finetune a lightweight GPT-2 model for translation using a next-token prediction loss. During inference, we do not have access to the ground-truth Marlin G-code, which would be needed to determine the cutoff lines for pair creation, so we instead split our Sailfish input into smaller chunks of fixed-length.
@@
Table 3 shows the results of performing finetuning on differently sized subsets of Slice-100K. We denote a GPT-2 model finetuned on $n$ shapes from our dataset as GPT-$2^{(n)}$. Figure 5 shows example renderings of shapes that have undergone flavor translation by our GPT-25 model. We also include an example G-code that has been translated in the Appendix (Figure 11). We find that even finetuning on minimal subsets of our dataset leads to significantly enhanced G-code translation abilities. Increasing the amount of training data beyond just five G-code shapes finetuning ceases to yield improvements. We attribute this to our preprocessing methods, which reduce the complex translation task to a simple local mapping, thereby reducing the amount of data needed for learning.
@@ 5 Conclusion
In this paper, we presented Slice-100K, the first large-scale, curated dataset of over 100,000 G-code files, along with their corresponding STL CAD files, renderings, and geometric properties. This dataset addresses a significant gap in the availability of comprehensive resources for additive manufacturing. We showcase our dataset explorer in Figure 6 and various modalities offered by Slice-100K in Figure 7. Our evaluation demonstrated the usefulness of Slice-100K in utilizing existing language models for tasks such as G-code debugging, geometric transformations, and comprehension. Additionally, we introduced a novel application of using Slice-100K for LLM-based G-code flavor translation, showcasing the potential of our dataset in advancing the field. We believe that Slice-100K will serve as a foundational resource for future innovations in manufacturing, paving the way for the development of domain-specific foundation models.
Limitations: Despite these advancements, Slice-100K has certain limitations. One major challenge is the difficulty in verifying the LVIS categories of the models. Additionally, all models in Slice-100K were sliced along the default Z-direction. This uniform slicing approach may limit the dataset’s applicability for research into multi-directional slicing techniques and their impact on manufacturing outcomes. Furthermore, we primarily focus on extrusion-based 3D printing among a plethora of additive manufacturing techniques. Addressing these limitations in future versions of the dataset will be crucial for further enhancing its utility and broadening its application scope.
@@ Acknowledgements
This work was supported by the National Science Foundation under grant CMMI-2347623/2347624. We would like to thank the NVIDIA Corporation for providing GPUs for academic research.
(a) Home screen
(b) Search manual
(c) Model view
Figure 6: Dataset explorer screenshots: (a) Home screen, (b) Search manual, and (c) Model view.
Figure 7: Various modalities offered by Slice-100K. First column - CAD models, Second column - G-code, Third column - Renderings, and Fourth and fifth columns - Geometric properties and LVIS categories.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Since the data generated involves fundamental science, the negative impact it has on society is of a similar nature to any fundamental manufacturing technology itself.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Material will be shared with reviewers.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See supplementary material.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See supplementary material.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See supplementary material.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [No]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See supplementary material.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See supplementary material.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Appendix
4:  $Lookup\leftarrow HashMap()$ $\triangleright$ Create hash index of contours in Layer B
5:  for $c\leftarrow 1$ to length(ContoursBContoursBdo
6:   for $i\leftarrow 1$ to length(ContoursB[c]ContoursB[c]do
@@
15:  $Mapping\leftarrow HashMap()$ $\triangleright$ Find bijection between layers
16:  for $c_{A}\leftarrow 1$ to length(ContoursAContoursAdo
@@
26:  for $i\leftarrow 1$ to length(ContoursBContoursBdo
@@
14:   if foundfound then $\triangleright$ Add matching pair of chunks to dataset
@@
19:     $end_{i}=end_{i}-1$ $\triangleright$ Could not find a line matching line $end_{i}$, try a smaller chunk